OpenStudy (anonymous):

Need help setting this up. H(x) = sqr root (x+6). Find the function h(7t-7), t > or = 1/7

OpenStudy (anonymous):

\[h(x)=\sqrt{x+6}\] \[h(7t-7)=\sqrt{7t-7+6}=\sqrt{7t-1}\]

OpenStudy (anonymous):

Ok so was the t greater than or equal to 1/7 sort of unnecessary info?

OpenStudy (anonymous):

\[if~7t-1<0,7t <1,t <\frac{ 1 }{ 7 }\] then we get negative value under the square root so h(7t-1) is imaginary. hence 7t-1 should be >= 0

OpenStudy (anonymous):

so the information given is not unnecessary.